sin a =12/13, sinb = 4/5. a lises in the 2nd quadrant and b lies in 1st quadrant. find tan(a+b)
 

Grade:11

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

sin A = 12/13
cos A = -5/13
tan A = -12/5

similarly, tan B = 4/3

now tan(A+B) = (tan A + tan B ) / ( 1- TanATanB )

plugin the values for the final answer

Regards
Vikas TU
3 years ago
Dear student
Ans would be 56/33
Hope this will help 
Good Luck,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,